OpenStudy (anonymous):

The product of (a − b)(a − b) is a^2 − b^2. My answer is always Am I right?

OpenStudy (agent0smith):

Expand it and find out if it's always equal... \[\Large (a − b)(a − b) = a^2 - ab - ab + b^2 = ...\]

OpenStudy (anonymous):

Its always, right?

OpenStudy (agent0smith):

Simplify what I gave, or at least try to, before you answer whether it's always or never.